Use the Pythagorean Theorem to find the distance to the nearest tenth, between F(9, 5) and G(-2, 2)
Answer by amoresroy(361)

Let d = the distance between points F and G
d^2 = (9- -2)^2 + (5-2)^2
d^2 = 121 + 9
d = 11.4